OpenStudy (anonymous):

If f(x) = 3x + 2 and g(x) = 2x – 2, what is (f – g)(x)? x + 4 x – 2 x 5x – 2 x – 4

OpenStudy (astrophysics):

This just means \[f(x)-g(x)\]

OpenStudy (astrophysics):

\[(f-g)(x) = 3x+2-(2x-2)\]

OpenStudy (anonymous):

I'm not sure what you mean?

OpenStudy (usukidoll):

it means the function of f(x) - the function of g(x)

OpenStudy (anonymous):

So you combine the two equations and that is the result?

OpenStudy (astrophysics):

\[(f-g)(x) = f(x)-g(x)\] subtract function g(x) from f(x).

OpenStudy (usukidoll):

yes... but please remember to distribute the negative for the right hand side of the equation

OpenStudy (usukidoll):

f(x) = 3x+2 and g(x) = 2x -2 so for (f-g)(x) which is the same as f(x) - g(x) it's the f(x) function minus the g(x) function so place your functions in here f(x) - g(x) then distribute that minus sign on the g(x) portion of the equation.

OpenStudy (usukidoll):

and we got to combine like terms afterwards.

OpenStudy (anonymous):

I think it's x + 4

OpenStudy (astrophysics):

Yeup!

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

thank you guys :)

OpenStudy (astrophysics):

Np